H andy-73-02 the bes e s by the Type here to search о не е
Answer #1

The Regression equation is ŷ = 7183 -0.294 x r=-0.247, naso when X=8 Y= 7.83 -(0.294) (8) = 7.83 - 2.352 û = 5,5 L Answer is